OpenStudy (anonymous):

Please help!! find all solutions of the following equation in the interval [0,2π) cos3x=-cos3x

OpenStudy (anonymous):

are you sure that you copied the question down correctly?

OpenStudy (anonymous):

oh theres a negative on the right side

OpenStudy (anonymous):

let m = 3x \[\cos(m) = -\cos(m)\] \[\cos(m) + \cos(m) = 0\] \[2\cos(m) = 0\] \[\cos(m) = 0\] \[m = \pi/2 * k \pi\] where k is an element of the integers \[3x = \pi/2 * k \pi\] \[x = \pi/6 * k \pi/3\]

OpenStudy (anonymous):

@alexandercpark i got that i just didnt know what the answer was... here are the choices: